# From CVE Entries to Verifiable Exploits: An Automated Multi-Agent Framework for Reproducing CVEs

### v1

@@ 4.4 Large Scale Evaluation
RQ3: At scale, CVE-Genie reproduced 428 / 841 CVEs, spanning 267 projects, 22 programming languages, and 141 CWE types, demonstrating broad applicability across ecosystems (see Table 4).
4.5 Importance of CVE-Genie’s Design
Ideal Case Knw Build Pre- Reqs Feed-back Critic Agents Single Agent
15 / 15 9 / 15 13 / 15 5 / 15 8 / 15 0 / 15
Table 5: Ablation study of CVE-Genie’s design.
Dataset. For this ablation study, we use the version of $D_{R}$ dataset which contains all essential information fields for 15 CVEs. As established earlier in Section 4.3, omitting these critical fields can affect accurate CVE reproduction. Therefore, in this ablation, we ensure all CVEs retain their complete critical information. This allows us to isolate the effect of changes to the design of CVE-Genie. If performance degrades when a specific change is made, we can attribute that drop solely to that change, since the CVE inputs remain otherwise fully informative.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
    Single Monolithic Agent: Combine all modular agents into a single agent with access to same tools, thereby evaluating the performance of a standalone LLM without CVE-Genie’s agentic design and structured guidance.
Results. In our ablation study (Table 5), we find that removing the Knowledge Builder drops reproduction to 9/15, as agents struggled with unstructured advisories and lost context, causing 30% more exploit failures. Eliminating the Pre-Requisite Developer had a mild impact (13/15), but overloaded the Setup Developer, which lead to 27% more tool-call limits being hit, particularly on complex repositories such as CVE-2025-32389 in Nameless, a website software for Minecraft servers. Feedback loops proved most critical: without them, performance collapsed to 5/15 (–67%). On the other hand, dropping critic agents cut results to 8/15 and raised false reproductions by 47%. Finally, collapsing into a single agent yielded 0/15 reproductions, even advanced standalone LLMs (o3) failed, underscoring the necessity of CVE-Genie’s modular, feedback-driven, critic-guided design.
RQ4: Standalone LLMs are currently unable to reproduce CVEs end-to-end. CVE-Genie’s modular design remains the most effective approach for enabling LLMs to perform this task comprehensively (Table 5).
@@ 5 Limitations and Future Work
Vulnerabilities that Involve UI Interactions. Currently, CVE-Genie only supports command-line interface (CLI) interactions. However, many web-based vulnerabilities require interaction with a graphical UI, making them difficult to trigger via CLI. Future work should explore integrating CVE-Genie with UI environments such as web browsers.
Multimodal CVE Knowledge Curation. Due to lack of image and video processing support, CVE-Genie struggles with CVEs where the PoC is provided as an image or video. Future work can address this limitation by integrating CVE-Genie with multimodal LLMs. Additionally, by leveraging external CVE resources through LLM-based deep search tools, beyond standard vulnerability databases, may also help improve reproduction success rates [38].
Over Critique. As shown in Figures 5(b), 6(b), and 7(b), all critics exhibit a lower TPR compared to their TNR. While this helps reduce false positives, it also results in missing some valid reproductions, particularly in cases where critics requested detailed setup, exploit, or verifier information that the developer could not sufficiently provide.
Cost. Running CVE-Genie costs approximately \$2.77 per CVE. While open-source models do not perform well and their large-scale deployment remains a challenge. To support further research, we release logs from all runs as a novel dataset to help the community fine-tune and improve local models’ performance for CVE reproduction.
@@ 6 Conclusion
In this paper we presented CVE-Genie, an automated framework designed to reproduce real-world vulnerabilities at scale. This has considerable benefit for downstream security research (see Appendix B), which suffers from a shortage of diverse high-quality data suitable for research and development of new tooling for automated vulnerability discovery and repair. CVE-Genie successfully reproduced around 51% (428) of 841 CVEs, with on average \$2.77 cost per CVE, across a vast variety of projects, CWEs, and programming languages. To the best of the authors’ knowledge, no other framework is capable of such comprehensive results at this scale and efficiency.
Ethical Considerations
The goal of our work is to support defenders and solution developers by accelerating vulnerability triage and tool development for vulnerability detection and patching. CVE-Genie, which does not create or determine new vulnerabilities, provides a powerful framework capable of providing recreations of CVE reports to provide up-to-date and reproducible CVE environments as benchmark test cases. These benchmarks can then be used to rigorously test and enhance vulnerability detection and patching systems, thereby understanding their performance and ideally accelerating their response capabilities. Following the beneficence criterion of the Menlo Report [14], we are confident that our system provides more benefits than potential harms to the security community and to society.
Open Science
We will open source CVE-Genie’s source code, logs for all experiments in Section 4 including all agent conversations, and our dataset of 428 reproduced CVEs.
Appendix A CVE-Genie Specifications
@@ Appendix B Applications of CVE-Genie
Vul.
detection
ML-based detection Training and testing data
@@
Vul.
triage
Root cause analysis Enrich inputs and enable more tools
@@
Vul.
Patching
Patch generation Benchmarks and testing environments
@@
Other
applications
LLM secure code generation Benchmarks and testing environment

### v2

@@ 4.4 Large Scale Evaluation
RQ4: At scale, CVE-Genie reproduced 428 / 841 CVEs, spanning 267 projects, 22 programming languages, and 141 CWEs, demonstrating broad applicability across ecosystems. Notably, 217 (51%) reproduced CVEs fall within the MITRE top 25 most dangerous CWEs, and 231 (54%) had no PoC available.
@@ 5 Limitations and Future Work
Vulnerabilities Involve UI Interactions. Currently, CVE-Genie only supports command-line interface (CLI) interactions. However, many web-based vulnerabilities require interaction with a graphical UI, making them difficult to trigger via CLI. Future work should explore integrating CVE-Genie with UI environments, e.g., web browsers.
Multimodal CVE Knowledge Curation. Due to a lack of image and video processing support, CVE-Genie struggles with CVEs where the PoC is provided as an image or video. Future work can address this limitation by integrating CVE-Genie with multimodal LLMs. Additionally, leveraging external CVE resources through LLM-based deep search tools, beyond standard vulnerability databases, may also help improve reproduction success rates [38].
Over Critique. As shown in Figures 5(b), 6(b), and 7(b), all critics exhibit a lower TPR compared to their TNR. While this helps reduce false positives, it also results in missing some valid reproductions, particularly in cases where critics requested detailed setup, exploit, or verifier that the developer could not sufficiently provide.
Cost. Running CVE-Genie costs approximately \$2.77 per CVE, moreover, open-source models currently do not perform well. Therefore, to support further research, we release logs from all reproduction runs as a novel dataset to help the community fine-tune and improve local models’ performance for CVE reproduction.
@@ 6 Conclusion
In this paper we presented CVE-Genie, an automated framework designed to reproduce real-world vulnerabilities at scale. This has considerable benefit for downstream security research (see Appendix F), which suffers from a shortage of diverse high-quality data suitable for research and development of new tooling for automated vulnerability discovery and repair. CVE-Genie successfully reproduced around 51% (428) of 841 CVEs, with on average \$2.77 cost per CVE, across a vast variety of projects, CWEs, and programming languages. To the best of the authors’ knowledge, no other framework is capable of such comprehensive results at this scale and efficiency.
Appendix A Acknowledgments
We gratefully acknowledge and thank Dave Aitel, Harold Nguyen, Xin Hu, and Ian Brelinsky (OpenAI), as well as Stijn Pletinckx and Lukas Dresel (University of California, Santa Barbara), for their valuable discussions and guidance that contributed to this work.
Appendix B Ethical Considerations
Safe Release of Reproduced CVEs’ Data. For all of our studies, we only include the CVEs that have a public patch available on the National Vulnerability Database (NVD). This measure prevents the attackers from using our reproduced CVEs to exploit the stakeholders’ systems, and also helps us to safely release the data of the CVEs reproduced in our experimental studies.
Release of CVE-Genie’s Codebase. The aim of releasing CVE-Genie’s codebase is to support defenders and solution developers by accelerating vulnerability triaging and tool development for vulnerability detection and patching. Although it is also possible for the attackers to leverage CVE-Genie to reproduce vulnerabilities in the real world, our aim is to level the playing field for defenders; specifically, to empower defenders to reproduce CVE reports to attain up-to-date and reproducible CVE environments as benchmark test cases. These benchmarks will then be used to rigorously test and enhance vulnerability detection and patching systems, thereby understanding their performance and ideally accelerating their response capabilities. Following the beneficence criterion of the Menlo Report [14], we strongly believe that our system provides more benefits than potential harms to the security community and to society.
Appendix C Open Science
CVE-Genie Framework and Dataset. https://github.com/BUseclab/cve-genie. This includes (1) source code of CVE-Genie and detailed instructions on how to run it, (2) reproduction runs of all reproduced CVEs with their agent conversations/logs, intermediate artifacts, and final exploit and verifier scripts, and (3) a web application to visualize the CVE reproduction runs.
Experiments Results. https://osf.io/dcej4/overview?view_only=81309bbc7e8c489abb9fb7e40bd1c0fb. This includes all reproduction runs with their agent conversations/logs, intermediate artifacts, and final scripts. (1) critics_evals.tar.gz contains results for selecting optimal LLMs and prompts (Section 3.5). (2) baseline.tar.gz contains results for the baseline assessment study (Section 4.1). (3) design_ablation.tar.gz contains results for CVE-Genie’s design ablations (Section 4.2). (4) robustness.tar.gz contains results for CVE-Genie’s robustness to loss of CVE context (Section 4.3). (5) large_scale.tar.gz contains results for multiple runs of the large-scale study with 841 CVEs (Section 4.4).
Appendix D Generative AI Usage
We used ChatGPT (OpenAI) to assist with LaTeX formatting, as well as to review and refine spelling, grammar, and sentence structure for improved clarity of the manuscript. Moreover, all AI-assisted outputs were manually reviewed, edited as needed, and explicitly approved by the authors before being incorporated into the paper.
Appendix E CVE-Genie Specifications
@@
Vul. detection ML-based detection Training and testing data
@@
Vul. triage Root cause analysis Enrich inputs and enable more tools
@@
Vul. Patching Patch generation Benchmarks and testing environments
@@
Other applications LLM secure code generation Benchmarks and testing environment
@@ Appendix F Applications of CVE-Genie
Appendix G Single Monolithic Agent Study Design
To evaluate a standalone LLM without CVE-Genie’s modular, agentic design and structured guidance, we collapsed all agents into a single monolithic agent with access to the full toolset (Appendix A), its entire conversation history, and an increased limit of 100 tool calls per attempt. The agent was provided with all CVE-related information directly in the user prompt, and the system prompt explicitly instructed it to follow a three-stage workflow (project setup, exploit generation, and verification), mirroring CVE-Genie’s high-level procedure. Unlike CVE-Genie, no automated critic or feedback loops were used; instead, we manually inspected execution traces and the generated exploit and verifier artifacts to determine whether reproduction was successful.
Appendix H Case Study for CVE-2024-5129
Overview. Versions of the lunary project prior to v1.2.8 expose a DELETE /v1/datasets/:id endpoint that lacks authorization checks and performs an unscoped deletion by dataset UUID. In the vulnerable release v1.2.7, the handler executes a SQL DELETE using only id, with no authentication and no restriction to the caller’s project, allowing unauthenticated deletion of arbitrary datasets by guessing or obtaining UUIDs.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-502 6 3 1 1 1
CWE-29 3 2 1 0 0
CWE-20 3 1 2 0 0
CWE-77 3 0 2 1 0
CWE-94 2 0 1 0 1
Table 9: Top CWEs reproduced in AI/ML Platform projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-79 99 34 37 10 18
CWE-284 29 12 8 2 7
CWE-22 21 7 6 3 5
CWE-200 19 5 3 2 9
CWE-918 16 4 5 1 6
CWE-89 15 5 6 1 3
CWE-400 15 6 3 1 5
CWE-20 12 4 3 1 4
CWE-285 9 5 2 2 0
Table 10: Top CWEs reproduced in Web App projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-79 33 9 6 12 6
CWE-1333 15 8 1 4 2
CWE-400 13 5 1 6 1
CWE-200 13 9 0 1 3
CWE-770 13 5 0 5 3
CWE-22 8 3 1 2 2
Table 11: Top CWEs reproduced in Lib/Framework projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-285 4 0 1 0 3
CWE-209 3 0 2 0 1
CWE-200 3 0 0 1 2
CWE-269 2 1 1 0 0
CWE-287 2 0 0 0 2
Table 12: Top CWEs reproduced in Cloud/DevOps projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
No-ID 7 0 0 5 2
CWE-122 3 0 0 3 0
CWE-400 2 0 0 1 1
CWE-416 2 0 0 0 2
CWE-863 2 0 0 0 2
Table 13: Top CWEs reproduced in OS/Runtime projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-79 7 0 6 0 1
CWE-122 4 0 0 1 3
CWE-416 4 0 0 2 2
CWE-22 2 0 2 0 0
CWE-20 2 0 1 0 1
Table 14: Top CWEs reproduced in Desktop App projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
No-ID 4 0 0 1 3
CWE-345 3 2 0 1 0
CWE-116 2 0 0 1 1
CWE-147 2 0 0 1 1
CWE-150 2 0 0 1 1
Table 15: Top CWEs reproduced in CLI Tool/Utility projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-670 3 1 1 0 1
CWE-200 2 0 2 0 0
CWE-400 1 0 1 0 0
Table 16: Top CWEs reproduced in Blockch/Crypto projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-532 2 0 0 1 1
CWE-261 1 0 0 1 0
CWE-327 1 0 0 1 0
Table 17: Top CWEs reproduced in MobileApp/SDK projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-200 4 1 0 1 2
CWE-416 3 0 0 1 2
CWE-670 2 1 0 0 1
CWE-639 2 1 0 0 1
Table 18: Top CWEs reproduced in Emb/Network projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-918 3 0 2 0 1
CWE-863 2 0 0 2 0
CWE-23 1 1 0 0 0
Table 19: Top CWEs reproduced in Security/Server projects.
Processor. The workflow starts with Data Processor resolving CVE-2024-5129’s affected versions and automatically downloading the latest vulnerable release (v1.2.7) from the upstream repository. Data Processor also extracts the CVE data, CWE information, patch commit, and bug bounty report for the CVE. The Knowledge Builderthen constructs a structured CVE knowledge base summarizing the vulnerability as missing authorization (CWE-862) and localizing the root cause to v1/datasets/index.ts, where the vulnerable delete handler deletes by dataset ID alone. The knowledge base also records the upstream fix (adding authorization middleware and scoping deletion by projectId), which becomes the ground truth for later exploit and verifier synthesis.
Builder. Given the vulnerable source tree and knowledge base, the Builder reconstructs a runnable environment. The Pre-Requisite Developer agent inspects repository structure and reads key configuration files (e.g., root package.json, backend package.json, backend .env.example, migration runner src/migrate.ts) to infer the required service stack: a Node.js/Koa backend backed by PostgreSQL and initialized via SQL migrations under packages/db.
The Setup Developer agent then provisions a real PostgreSQL 14 service inside the VM, configures credentials, and generates a valid .env file pointing DATABASE_URL to the local database instance. All Node.js dependencies are installed via npm install, and database migrations are executed using the project’s native migration runner (npm run migrate:db), resulting in a fully initialized schema. Finally, the backend is launched using npm run dev and confirmed operational by observing successful database connectivity and a listening API server on port 3333. The Setup Critic agent analyzes the execution logs of Setup Developer and approves the build.
After environment reconstruction, the Setup Critic evaluates the entire setup. The critic confirms that (i) the vulnerable code path is present and unchanged, (ii) a real PostgreSQL-backed Lunary backend is running, (iii) no dummy services or mocked endpoints were introduced, and (iv) the vulnerable DELETE /v1/datasets/:id endpoint is reachable and responds with HTTP 200 to unauthenticated requests. Based on these checks, the critic accepts the project build as a genuine and exploitable reproduction target.
Exploiter. The Exploit Developer agent begins by performing targeted code inspection to ground the exploit in the real implementation. It reads the vulnerable dataset router source file and confirms that the datasets.delete("/:id") handler executes an unscoped SQL deletion without any authentication or authorization middleware. It also inspects backend configuration files to verify the correct runtime entry points and dependencies. To validate exploitability in practice, the exploiter launches the backend and issues unauthenticated DELETE requests against /v1/datasets/<uuid> using both ad-hoc Python test scripts and raw HTTP calls. A randomly generated UUID is sufficient to elicit an unconditional HTTP 200 response, confirming that no permission checks or project scoping are enforced. Based on this confirmation, the exploiter generates a standalone PoC exploit.py script. The script accepts a dataset UUID (and an optional base URL), sends an unauthenticated DELETE request to the vulnerable endpoint, and prints the HTTP status code and response body. The exploit is intentionally minimal, containing no setup logic or auxiliary assumptions beyond network access to the API.
The generated exploit artifact is then evaluated by the Exploit Criticagent. The critic verifies that the exploit (i) targets the correct vulnerable endpoint, (ii) directly triggers the missing-authorization flaw described in the CVE knowledge base, (iii) does not rely on environmental shortcuts or fabricated behavior, and (iv) has been empirically validated against the running vulnerable service. After confirming these properties, the critic accepts the exploit as a correct and faithful PoC.
CTF Verifier. To automatically confirm successful exploitation of CVE-2024-5129, the Verifier Developer agent synthesizes a CTF-style verifier that executes the exploit and validates its effect against the reconstructed Lunary backend. The verifier follows a fixed the following three-phase structure.
  1. 1.
    In the pre-setup phase, it inserts a minimal but valid set of rows, i.e., organization, account, project, and dataset, directly into the PostgreSQL database used by the running backend. A fresh dataset UUID is generated and recorded, ensuring that a concrete deletion target exists prior to exploitation.
  2. 2.
    In the exploit phase, the verifier invokes the previously generated exploit.py script verbatim via subprocess.run, passing the dataset UUID and the base URL http://localhost:3333. The verifier itself performs no HTTP requests, ensuring that the exploit remains the sole trigger of the vulnerability.
  3. 3.
    In the post-exploit phase, the verifier re-queries the database to determine whether the dataset row with UUID still exists. If the row has been deleted, the verifier prints the flag 3xploit66full; otherwise, it reports failure. Database access is implemented robustly, using psycopg2 when available and falling back to the psql client if necessary.
During verification, earlier verifier iterations that relied on mocked or stubbed HTTP servers were rejected by the Verifier Critic agent. The final accepted verifier operates exclusively on the real Node.js / Koa backend and PostgreSQL instance and performs a direct state-based validation. The Flag Checker successfully executes the verifier, observes flag emission, and the Verifier Critic confirms the verifier as a correct and faithful reproduction of the CVE.
Storing Artifacts and Metadata. After the successful verification, CVE-Genie stores the VM snapshot with running Lunary server. Moreover, it stores exploit and verifier scripts, agent logs, and metadata of the reproduction run. Overall, this reproduction run of CVE-2024-5129 cost \$1.68 and 20 minutes. See artifacts here  [^5] https://github.com/BUseclab/cve-genie/tree/main/results/reproduced_cves/CVE-2024-5129 in open-source dataset.